# Electric Vehicle Charging Stations Placement Optimization in Vietnam Using Mixed-Integer Nonlinear Programming Model

### v1

@@ Abstract
Vietnam is viewed as one of the promising markets for electric vehicles (EVs), especially automobiles when it is predicted to reach 1 million in 2028 and 3.5 million in 2040. However, the lack of charging station infrastructure has hindered the growth rate of EVs in this country. This study aims to propose an optimization model using Mixed-Integer Nonlinear Programming (MINLP) to implement an optimal location strategy for EVs charging stations in Ho Chi Minh (HCM) City. The problem is solved by a solver named Gurobi and using the Brand-and-Cut method. There are 2 perspectives including Charging Station Operators and EV users. In addition, 7 kinds of costs considered include installation cost, land rental cost, maintenance cost, operational cost, charging cost, waiting cost, and traveling cost. From 1509 Point of Interest and 199 residential areas, 134 POIs were chosen with 923 charging stations including 592 Level-2 chargers and 331 Level-3 chargers to fully satisfy the customer demand. Furthermore, the effectiveness of the proposed model is proved by a minor MIP Gap and running in a short time with full feasibility.
Index Terms:
@@
I Introduction
In the new global economy, environmental issue has become an important worldwide problem. Transport sector accounts for magnification of this, with million of gasoline-powered vehicles on the roads emitting carbon continuously, carbon emissions persist at a high rate. Besides, gasoline-powered car also create cacophonous pollution from internal combustion engine [1]. In order to alleviate these vulnerabilities and promote a more sustainable economy, one solution is to switch from gasoline-based vehicles to using green technology vehicles. For Vietnam, these advantages align closely with the country’s sustainability objectives, helping move it closer to its Net Zero goal by 2050.
@@ I Introduction
The contributions of this paper can be summarized in three key aspects.
  • Firstly, this study employs an extended mathematical model to simultaneously explore the perspectives of both station operators and users, proposing solutions that balance the interests of providers and consumers while addressing gaps in prior research.
  • Secondly, it is one of the few domestic studies to evaluate the economic aspects of stations. By applying the Mixed-Integer Nonlinear Programming (MINLP) model - to a focused area (determining the optimal number of stations - ##\##stations) and within a new survey scope (HCM City), this study adds a novel dimension.
  • Finally, the research offers strategic and up-to-date insights that are directly applicable to business practice. The arguments derived from this study can provide decision-makers in HCM City with valuable information, improve strategic planning for station development, help attract and serve the right audience in the right locations.
The analysis identifies Thu Duc City as having the highest POIs (19.02%) and population (13.06%), highlighting its strong potential for EV charging infrastructure. High demand is also observed in District 12, District 7, and Binh Thanh District, while suburban areas like Can Gio District and Nha Be District require infrastructure expansion despite lower demand. Newly installed stations are concentrated in central urban and underserved suburban areas, with Level-2 stations generally outnumbering Level-3 stations, except in Binh Chanh District and Thu Duc City. Charging costs dominate total expenses (35.47%), followed by operational costs (28.83%) and travel costs (27.99%).
The first section provides an overview of the research. The next two parts present the theoretical and empirical basis for the study, along with the models and algorithms used. The fourth section describes the research methodology, detailing the data collection methods, data sources, and data processing techniques. Next, the input data and evaluation criteria for assessing the model is presented. The model is then evaluated based on the metrics described in the methodology section. The final section concludes the study, drawing insights from the results and offering practical recommendations. Limitations of the research are also identified, providing directions for future studies in related areas.
II Literature Review
@@
TABLE I: Summary of previous research
Reference Objective function Model Metrics Dataset
@@
[4] Minimizing voltage fluctuations MINLP Time Consumption Networking and Emerging Optimization and IEEE Power and Energy Socierty
@@
[15] Maximum Coverage Location Problem (MCLP) LP (PuLP with GLPK package) Wake County Government GIS, Google Maps
@@
A large number of research have proliferated and can be summarized into five perspectives: DNOs, environment approach, CSOs, EV users, and consumer behavior. These research directions are summarized in the Figure 1 and Table  I. Furthermore, five approaches are elaborated below.
@@
Research related to EV users, the owners that responsible for recharging their vehicles at these stations. Since most stations have listed prices, the primary concern for EV owners is the distance they must travel to reach a station. Studies on problems like the Maximum Coverage Location Problem (MCLP) [15] and the knapsack problem [12] focus heavily on this aspect. Additionally, in countries that support and implement Vehicle-To-Grid (V2G) technology - which enables energy to flow back from EVs to the grid - research aims to optimize the timing and battery percentage that station users adjust to assist the grid during emergencies or manage charging and discharging during off-peak and peak hours [16], [17]. The overarching objective of these studies is to strategically place stations in locations that minimize travel costs while effectively meeting the charging demands of surrounding areas. Furthermore, optimizing the costs associated with stations - including charging costs [18], waiting time costs [19], and travel costs [20] - is a significant focus in this research field.
@@
Studies on consumer behavior focus on analyzing the factors that directly motivate consumers to consider purchasing EVs or to transition from traditional gasoline cars to this novel mode of transportation.[21] expanded the Technology Acceptance Model (TAM) to explore how perceived risks - namely financial, functional, and social risks - impact purchasing decisions. Meanwhile, other studies have highlighted various factors such as perceived benefits, ease of use, facilitating conditions, cost perception, usage habits, and environmental awareness, perceived usefulness, ease of use, social influence, social barriers, charging infrastructure, and environmental awareness in shaping consumer decisions. In summary, these studies collectively provide valuable insights into the motivations and hesitations influencing the shift towards EVs.
@@
In Vietnam, research has predominantly focused on technical issues, such as electrical load, reliability, batteries, and consumer behavior in purchasing decisions. Conversely, economic aspects related to EVs remain unexplored, presenting a new area for further investigation. Therefore, focusing on DNOs, CSOs, and EV users offers promising research directions. Regarding the placement of stations to optimize the economic benefits for grid operators (DNOs), according to [25]: “The State (Vietnam Government) holds a monopoly over national electric system regulation” . This means that in Vietnam, the Government retains exclusive control over the electricity system, particularly transmission. Consequently, studies focused on grid operation are challenging to apply within the Vietnamese context. Thus, the perspective of CSOs and EV users serves as the primary direction for this research. Building on the Vietnamese context and insights from international studies, this research aims to incorporate previous findings and address existing research gaps.
III Backgrounds
TABLE II: Notation and definitions used in the MINLP model
@@ III Backgrounds
The problem of optimizing EV stations involves numerous variables, classifying it among highly complex (NP-hard) problems. To date, various methods have been employed to address these optimization challenges, including mathematical optimization and model-based programming techniques. The notation, as presented in Table II, is presented throughout the MINLP model.
III-A Mixed-Integer Nonlinear Programming Model
In this research, we choose the Mixed-Integer Nonlinear Programming model (MINLP) which is widely used in recent studies [26, 27, 28], as it integrates linear, nonlinear, and mixed-integer programming. This technique can handle both continuous and discrete variables along with nonlinear functions, providing reliable solutions for multi-objective problems. The MINLP model is particularly suitable for handling complex problems that involve nonlinearity and integer variables, often requiring extensive search trees for solutions [29]. Specifically, the target is that minimizing objective function $f(x)$ with constraint $c(x)$:
@@ IV-B Objective function
across $n$ locations when installing and operating a station from CSOs and EV users perspectives, taking into account 7 cost components associated with the deployment and operation of stations. Table III summarizes the definitions and meanings of the cost components denoted in (2).
TABLE III: Summary of 7 different components
@@
$L_{i}$ Land rental cost The expense of securing space for the station daily. The formula construction is presented detail in Appendix. A. $\frac{l_{i}}{14600}(x_{2,i}+x_{3,i})$
@@ IV-C Constraints
V Experiments
Figure 3: Experiments schedule with the crawled dataset from Section. V, passed through Gurobi solver.
V-A Dataset
Fig. 3 illustrates the data processing and optimization workflow for determining electric vehicle station locations using the MINLP model. Raw data is collected from 3 sources including OpenStreetMap, Google Map API, and the HCM City Open Data Portal, then processed through geocoding to convert it into coordinates (latitude and longitude). The input data is organized into two tables: Point of Interest (POI) (including ID, coordinates, and land cost) and RP (including ID, coordinates, and population). The MINLP model utilizes this data to define the objective function and constraints, which are then solved using the Gurobi Solver. The process concludes with an optimal solution that determines the locations and quantities of stations to best meet the model’s requirements.
In demand modeling, the population is often used as a representative of that region, [26] divided the study area into smaller regions, assuming that residents concentrate at the region’s center, in which the area is divided into regions based on estimated EV density and charging demand, with distributed stations ensuring access for all EV users [26]. Thus, this study uses population data by ward to represent demand in each area, with each ward centered around its local administrative office, it is also called Residential Point (RP). As of December 31, 2022, the city has 22 district-level administrative units (including 5 rural districts and 1 municipal city) and 249 wards [40], the POIs and RPs are distributed not equally between districts, as shown in Fig. 6. Table IV summarizes the 1,509 POIs used as input for the research model, categorized into six main groups: parking lots, supermarkets/shopping malls, apartments/office buildings, universities/colleges, hotels, and gas stations. Among these, apartments/office buildings have the highest number (521 locations, 34.5%), followed by gas stations (514 locations, 34%). Places such as hotels (61 locations, 4%) and parking lots (94 locations, 6.2%) account for smaller proportions, reflecting an uneven distribution. The diverse dataset, ranging from public spaces to commercial areas and existing infrastructure, highlights that the model has carefully considered various aspects to optimize the placement of stations.
TABLE IV: Statistics of POI Data
Category Label Total Number Source Supermarket, Shopping Mall market, supermarket, grocery store, hypermarket, mall 128 Google Maps Gas Station pvoil, gas station 514 [41] Parking Lot parking 94 Google Maps Apartment, Office Building, Tower apartment, office building, office space, tower, business tower 521 Google Maps, [42] University/College university, college, higher education institution, academic institution 191 Google Maps Hotel hotel, accommodation, lodge, resort 61 Google Maps
V-B Metrics
@@ VII Conclusion
Currently, the establishment of standards for electric stations in Vietnam is being developed. In the future, the construction of stations will need to comply with these regulatory standards, thus not all potential POIs will meet the conditions for installation. Therefore, future studies should consider these standards and regulations when selecting locations for stations, ensuring that models are up-to-date. In subsequent research, expanding the scope to survey additional provinces across the country would provide a more inclusive reflection of the findings. Additionally, if time and resources allow, the research could be further enriched by gathering insights from experts in relevant fields, thereby enhancing the comprehensiveness and reliability of the results.
@@
The codes and data used for this study are available at https://github.com/tulliavu/EVCSVN.
@@ Acknowledgment
Appendix A Land rental cost
The land rental cost represents an opportunity cost that landowners or CSOs must consider when deciding to set up and operate a station at a specific location. Land rental costs are calculated based on the product of the minimum area required to accommodate an EV, the land rental price and the number of charging ports [12]. In Vietnam, the National Technical Regulation on Construction Planning requires a minimum of 25 $m^{2}$ per parking spot on the ground level [46]. On January 16, 2020, the HCM City People’s Committee issued [47], specifying land values for streets across the city from 2020 to 2024. After that, a draft adjustment for [47] was released on July 19, 2024, to update the land values, ensuring they reflect current market conditions. The latest land pricing data from this draft was used as input for the algorithm in this study, as it provides the most current land valuation in line with present-day circumstances.

### v2

@@ Abstract
Vietnam is viewed as one of the promising markets for electric vehicles (EVs), especially automobiles, when it is predicted to reach 1 million in 2028 and 3.5 million in 2040. However, the lack of charging station infrastructure has hindered the growth rate of EVs in this country. This study aims to propose an optimization model using Mixed-Integer Nonlinear Programming to implement an optimal location strategy for EVs charging stations in Ho Chi Minh City. The problem is solved by Gurobi using the Brand-and-Cut method. There are two perspectives, including Charging Station Operators and EV users. In addition, 7 kinds of costs are considered. From 1509 Point of Interest and 199 residential areas, 134 POIs were chosen with 923 charging stations to fully satisfy the customer demand. Furthermore, the effectiveness of the proposed model is proved by a minor MIP Gap and running in a short time with full feasibility.
keywords:
@@
articletype: ARTICLE TEMPLATE
1 Introduction
In the new global economy, environmental issues have become an important worldwide problem. Transport sector accounts for magnification of this, with millions of gasoline-powered vehicles on the roads emitting carbon continuously, carbon emissions persist at a high rate. Besides, gasoline-powered cars also create cacophonous pollution from the internal combustion engine [1]. In order to alleviate these vulnerabilities and promote a more sustainable economy, one solution is to switch from gasoline-based vehicles to using green technology vehicles. For Vietnam, these advantages align closely with the country’s sustainability objectives, helping move it closer to its Net Zero goal by 2050.
@@ 1 Introduction
The contributions of this paper can be summarized in two key aspects.
  • This study simultaneously explores the perspectives of both station operators and users, proposing solutions that balance the interests of providers and consumers.
  • It evaluates the economic aspects of stations within a new survey scope (HCM City).
As a result, the analysis identifies Thu Duc City as having the highest POIs (19.02%) and population (13.06%), highlighting its strong potential for EV charging infrastructure. High demand is also observed in District 12, District 7, and Binh Thanh District, while suburban areas like Can Gio District and Nha Be District require infrastructure expansion despite lower demand. Newly installed stations are concentrated in central urban and underserved suburban areas, with Level-2 stations generally outnumbering Level-3 stations, except in Binh Chanh District and Thu Duc City. Charging costs dominate total expenses (35.47%), followed by operational costs (28.83%) and travel costs (27.99%).
2 Literature Review
@@
Table 1: Summary of previous research
Ref Objective function Model Metrics Dataset
@@
[4] Minimizing power loss, network capacity and voltage regulation HGAMSSA Convergence process IEEE 33 bus distribution network
@@
[15] Maximum Coverage Location Problem (MCLP) LP (PuLP with GLPK package) Wake County Government GIS, Google Maps None
@@
A large number of research has proliferated and can be summarized into five perspectives: DNOs, environment approach, CSOs, EV users, and consumer behavior. These research directions are summarized in the Figure 1 and Table  1. Furthermore, five approaches are elaborated below.
@@
Research related to EV users, the owners who are responsible for recharging their vehicles at these stations. Since most stations have listed prices, the primary concern for EV owners is the distance they must travel to reach a station. Studies on problems like the Maximum Coverage Location Problem (MCLP) [15]. Additionally, in countries that support and implement Vehicle-To-Grid (V2G) technology - which enables energy to flow back from EVs to the grid - research aims to optimize the timing and battery percentage that station users adjust to assist the grid during emergencies or manage charging and discharging during off-peak and peak hours [16], [17]. The overarching objective of these studies is to strategically place stations in locations that minimize travel costs while effectively meeting the charging demands of surrounding areas. Furthermore, optimizing the costs associated with stations - including charging costs [18], waiting time costs [19], and travel costs [20] - is a significant focus in this research field.
@@
Studies on consumer behavior focus on analyzing the factors that directly motivate consumers to consider purchasing EVs or to transition from traditional gasoline cars to this novel mode of transportation.[21] expanded the Technology Acceptance Model (TAM) to explore how perceived risks - namely financial, functional, and social risks - impact purchasing decisions. Meanwhile, other studies have highlighted various factors such as perceived benefits, ease of use, facilitating conditions, cost perception, usage habits, and environmental awareness, perceived usefulness, ease of use, social influence, social barriers, charging infrastructure, and environmental awareness in shaping consumer decisions. In summary, those studies regarding consumer behavior provide insights into the motivation and factors influencing customer decisions when shifting towards EVs.
@@
In Vietnam, since the electric vehicle industry has only recently emerged, related research has primarily focused on technical issues such as electrical load, reliability, batteries, and consumer behavior in purchasing decisions. Conversely, economic aspects related to EVs remain unexplored, presenting a new area for further investigation. Therefore, focusing on DNOs, CSOs, and EV users offers promising research directions. Regarding the placement of charging stations to optimize economic benefits for grid operators (DNOs), according to [25]: “The State (Vietnam Government) holds a monopoly over national electric system regulation.” This means that in Vietnam, the Government retains exclusive control over the electricity system, particularly transmission, making studies focused on grid operation challenging to apply within the Vietnamese context. At the same time, existing studies on EV charging station placement offer diverse perspectives; however, they are often based on simulated data [4] or tailored to specific regions such as the UK [12] and China [13]. These models may not be directly applicable to Ho Chi Minh City, given its unique geographical characteristics and public policy framework. Therefore, this research adapts existing methodologies to better align with the local context, ensuring a more practical and region-specific approach. Building on the Vietnamese context and insights from international studies, this research aims to incorporate previous findings and address existing research gaps.
3 Backgrounds
Table 2: Notation and definitions used in the MINLP model
@@ 3 Backgrounds
Ho Chi Minh City has been selected as the ideal location for developing an optimized electric vehicle (EV) charging station model due to its large population and the increasing demand for EVs. Despite covering only 0.63% of the country’s total area, with a land size of 2,095 km², the city had a population of over 9 million as of December 2022, accounting for nearly 10% of Vietnam’s total population [26]. This makes it not only the most populous city in Vietnam but also the country’s economic hub, characterized by a dynamic urban lifestyle and a high volume of traffic. Given these factors, Ho Chi Minh City presents a compelling case for implementing an optimized EV charging station model to support sustainable urban mobility.
The problem of optimizing EV stations involves numerous variables, classifying it among highly complex (NP-hard) problems. To date, various methods have been employed to address these optimization challenges, including mathematical optimization and model-based programming techniques. The notation, as presented in Table 2, is presented throughout the MINLP model.
3.1 Mixed-Integer Nonlinear Programming Model
In this research, we choose the Mixed-Integer Nonlinear Programming model (MINLP) which is widely used in recent studies [27, 28, 29], as it integrates linear, nonlinear, and mixed-integer programming. This technique can handle both continuous and discrete variables along with nonlinear functions, providing reliable solutions for multi-objective problems. The MINLP model is particularly suitable for handling complex problems that involve nonlinearity and integer variables, often requiring extensive search trees for solutions. Specifically, the target is that minimizing objective function $f(x)$ with constraint $c(x)$:
@@ 4.2 Objective function
across $n$ locations when installing and operating a station from CSOs and EV users perspectives, taking into account 7 cost components associated with the deployment and operation of stations. Table 3 summarizes the definitions and meanings of the cost components denoted in (2).
Table 3: Summary of 7 different components
@@
$L_{i}$ Land rental cost The expense of securing space for the station daily. $\frac{l_{i}}{14600}(x_{2,i}+x_{3,i})$
@@ 4.3 Constraints
5 Experiments
Figure 3: Experiments schedule with the crawled dataset from Section. 5, passed through Gurobi solver.
5.1 Dataset
Fig. 3 illustrates the data processing and optimization workflow for determining electric vehicle station locations using the MINLP model. Raw data is collected from 3 sources, including OpenStreetMap, Google Map API, and the HCM City Open Data Portal, then processed through geocoding to convert it into coordinates (latitude and longitude). The input data is organized into two tables: Point of Interest (POI) (including ID, coordinates, and land cost) and RP (including ID, coordinates, and population). The MINLP model utilizes this data to define the objective function and constraints, which are then solved using the Gurobi Solver. The process concludes with an optimal solution that determines the locations and quantities of stations to best meet the model’s requirements.
In demand modeling, the population is often used as a representative of that region, [27] divided the study area into smaller regions, assuming that residents concentrate at the region’s center, in which the area is divided into regions based on estimated EV density and charging demand, with distributed stations ensuring access for all EV users [27]. Thus, this study uses population data by ward to represent demand in each area, with each ward centered around its local administrative office, it is also called Residential Point (RP). As of December 31, 2022, the city has 22 district-level administrative units (including 5 rural districts and 1 municipal city) and 249 wards [26], the POIs and RPs are distributed not equally between districts, as shown in Fig. 6. Table 4 summarizes the 1,509 POIs used as input for the research model, categorized into six main groups: parking lots, supermarkets/shopping malls, apartments/office buildings, universities/colleges, hotels, and gas stations. Among these, apartments/office buildings have the highest number (521 locations, 34.5%), followed by gas stations (514 locations, 34%). Places such as hotels (61 locations, 4%) and parking lots (94 locations, 6.2%) account for smaller proportions, reflecting an uneven distribution. The diverse dataset, ranging from public spaces to commercial areas and existing infrastructure, highlights that the model has carefully considered various aspects to optimize the placement of stations.
Table 4: Statistics of POI Data
Category Label Total Number Source Supermarket, Shopping Mall market, supermarket, grocery store, hypermarket, mall 128 Google Maps Gas Station pvoil, gas station 514 [40] Parking Lot parking 94 Google Maps Apartment, Office Building, Tower apartment, office building, office space, tower, business tower 521 Google Maps, [41] University, College university, college, higher education institution, academic institution 191 Google Maps Hotel hotel, accommodation, lodge, resort 61 Google Maps
5.2 Metrics
@@ 7 Conclusion
Currently, Vietnam is in the process of developing regulatory standards for electric vehicle charging stations. In the future, station construction will need to comply with these standards, meaning that not all potential POIs will be eligible for installation. Therefore, future studies should incorporate these evolving regulations when selecting station locations to ensure models remain relevant and applicable. Additionally, expanding the scope of research to include multiple provinces across Vietnam would provide a more comprehensive and representative analysis. Beyond geographical considerations, future research should also integrate dynamic EV usage patterns and user preferences, as this study primarily relies on static data. Moreover, sensitivity analysis using different algorithms would help assess model robustness and computational complexity, enhancing its applicability in diverse scenarios. Lastly, if time and resources permit, incorporating expert insights from relevant fields could further refine the study, improving both the depth and reliability of the findings.
@@
The data that support the findings of this study are openly available at https://github.com/tulliavu/EVCSVN.